the sum of is three numbers is 36.the second is five more than the first, and the third is four less than three times the first. find the numbers.
Answer by ptth176(50)

the first number=x
the second number=y
the third number=z
the sum of is three numbers is 36=>x+y+z=36(1)
the second is five more than the first=>y=x+5(2)
the third is four less than three times the first=>z=3x-4(3)
(1)(2)(3)=>x+x+5+3x-4=36
=>5x=35
=>x=7
(2)=>y=7+5=12
(3)=>z=3(7)-4=17